# A Foundational Theory for Decentralized Sensory Learning

### v1

@@ 2 Conclusion
At the start of this text, we provided a theory that explained how the first unicellular life forms could survive by propagating problem signals from various cellular sensors across the cell to solve problems in order of priority.
@@
After that, we discussed how evolution could pressure the development of electrical signaling to allow the same problem propagation used by unicellular and multicellular organisms to function across larger distances and diverse sensors throughout a body.
@@
After that, we discussed how our theoretical framework can act as a unifying principle to support many experimental neuroscientific observations, and finally we also discussed why these sensory signals are the only necessary value signals for biological life and the brain to function and learn.

### v2

@@ 2 Conclusion
We provided a theory that explained how the first unicellular life forms could survive by propagating problem signals from various cellular sensors across the cell to solve problems in order of priority.
@@
After that, we discussed how evolution could pressure the development of electrical signaling to allow the same problem propagation used by unicellular and multicellular organisms to function across larger distances and diverse sensors throughout a body - the emergence of the nervous system.
@@
Furthermore, we presented a bare bones implementation based on the conjectures that was able to solve a classic control problem and demonstrated how other implementations of the conjectures have been applied to solve the control of quadruped and biped systems in real-world and in simulation, respectively.
We discussed how our theoretical framework can act as a unifying principle to support many experimental neuroscientific observations, and also why these sensory signals are the only necessary value signals for biological life and the brain to function and learn, as well as how this differs from modern artificial intelligence.
@@ 4 Conflict Of Interest
Supplementary Figures
Supplementary Figure 1: Grand average synaptic weight over normalized time for the simple cell system. Less saturated blue indicates ±2 standard deviations (s.d.s) for the difference between timesteps.
Supplementary Figure 2: Grand average synaptic weights over normalized time for the larger cell system, shown for connections from the Input layer (red), from and within the Recurrent layer (green), and all of the connections (blue). Less saturated colors indicate ±2 s.d.s for the differences between timesteps.
Supplementary Figure 3: Matrices of the relative connectivity strengths (calibrations in the color bar) between individual sensors and motor neurons (MN) at the end of the simulation for the 10 different random network initializations for the simple cell system.
Supplementary Figure 4: Matrices of the relative connectivity strengths (calibrations in the color bar) between individual sensors and intermediary neurons (‘recurrent’), within the intermediary layer, and from the intermediary layer to the motor neurons (MN) at the end of the simulation for the 5 different random network initializations for the larger cell system. The relative simplicity of the problem context, the inverted pendulum, is reflected in the comparatively sparse utilization of the recurrent-recurrent connectivity implying a low utility of latent space computation while simultaneously arriving at a stable problem solving solution.